# Bell Correlations via Constrained Colliders – a Path from W to V

### v1

@@ Abstract
In previous work with Ken Wharton, I have proposed that Bell correlations are a special sort of selection artefact, explained by a combination of (i) collider bias and (ii) a boundary constraint on the collider variable. This hypothesis requires no direct causal influence outside lightcones, and may hence offer a new way to reconcile Bell nonlocality and relativity. This piece outlines a new argument for the proposal. It explains how it is valid for a special class of (‘-shaped’) Bell experiments involving delayed-choice entanglement swapping, and argues that it can be extended to the general (‘-shaped’) case.
@@ 2.3 Collider bias in DCES versions of
Several authors have noted that in the delayed-choice version of the -shaped experiment, Bell correlations in the postselected ensemble are selection artefacts [11, 9, 10, 2, 28, 38]. A mark of this, as we have just seen, is that these correlations do not support the counterfactuals characteristic of Bell nonlocality. A change in Alice’s setting may produce a change in the result of the intermediate measurement M, thus removing the case from the postselected ensemble. There is hence no need for it make a difference to Bob’s outcome . In Bell experiments of this kind, then, the Bell correlations themselves are counterfactually fragile.
@@ 2.5.1 A delayed-choice -shaped experiment?
Note, however, that if the setting choices at A and B can make a difference to the outcome at D, then they are also making a difference at C. The case then involves retrocausality, as shown in red in Figure 4. Some readers will balk at this possibility. Invoking the words of John Wheeler, perhaps, they may object that it allows that ‘present choice influences past dynamics, in contravention of every principle of causality’ [35, p. 41]. It would take us too far afield to challenge Wheeler’s instincts on this matter.666Ken Wharton and I have done so at length elsewhere [23, 26, 27, 34]. But for readers wedded to those instincts, we need a more powerful intuition pump.
@@ 2.5.2 The constraint-free regime
In CFR, then – with these caveats about what we mean – we have no reason to exclude the possibility that settings choices and ‘influence’ the initial state at C, as shown in red in Figure 5. The initial state thus becomes a collider, and Bell correlations can now be explained as collider artefacts, just as in the delayed-choice -shaped case.888Some readers may be troubled by the presumption that there could be experiments at all in CFR. But here we are on well-trodden ground. A community familiar with Boltzmann Brains should not be troubled by a few simple pieces of quantum optics. Such things will arise by chance in a sufficiently expansive CFR, and we may discuss their characteristic statistics.

### v2

# W as the Edge of a Wedge: Bell Correlations via Constrained Colliders

@@ Abstract
In previous work with Ken Wharton, it was proposed that Bell correlations are a special sort of selection artefact, explained by a combination of (i) collider bias and (ii) a boundary constraint on the collider variable. This requires no direct causal influence outside lightcones, and may hence offer a new way to reconcile Bell nonlocality and relativity. This piece outlines a new argument for the proposal. It explains how it is valid for a special class of (‘-shaped’) Bell experiments involving delayed-choice entanglement swapping, and argues that it can be extended to the general (‘-shaped’) case.
@@ 2.3 Collider bias in DCES versions of
Several authors have noted that in the delayed-choice version of the -shaped experiment, Bell correlations in the postselected ensemble are selection artefacts [11, 9, 10, 2, 29, 19]. A mark of this, as we have just seen, is that these correlations need not support the counterfactuals characteristic of Bell nonlocality. A change in Alice’s setting may produce a change in the result of the intermediate measurement M, thus removing the case from the postselected ensemble. There is hence no need for it make a difference to Bob’s outcome . In Bell experiments of this kind, then, the Bell correlations themselves may be counterfactually fragile.
@@ 2.5.1 A delayed-choice -shaped experiment?
Note, however, that if the setting choices at A and B can make a difference to the outcome at D, then they are also making a difference at C. The case then involves retrocausality, as shown in red in Figure 4. Some readers will balk at this possibility. Invoking the words of John Wheeler, perhaps, they may object that it allows that ‘present choice influences past dynamics, in contravention of every principle of causality’ [36, p. 41]. It would take us too far afield to challenge Wheeler’s instincts on this matter.666Ken Wharton and I have done so at length elsewhere [24, 27, 28, 35]. In the present context, we may be overstating the relevance of Wheeler’s objection. Wheeler is discussing a delayed-choice experiment, and the full passage runs like this: Does this result mean the present choice influences past dynamics, in contravention of every principle of causality? Or does it mean, calculate pedantically and don’t ask questions? Neither; the lesson presents itself rather as this, that the past has no existence except as it is recorded in the present. Wheeler’s own option may well be sufficient for our present purposes. If in Figure 4 the measurement settings at A and B can make a difference to the past recorded at D, that seems good enough. But for readers wedded to those instincts, we need a more powerful intuition pump.
@@ 2.5.2 The constraint-free regime
In CFR, then – with these caveats about what we mean – we have no reason to exclude the possibility that settings choices and ‘influence’ the initial state at C, as shown in olive green in Figure 5. (We have changed the colour of the arrows to remind readers that causality means something different than in Figure 4.) The initial state thus becomes a collider, and Bell correlations can now be explained as collider artefacts, just as in the delayed-choice -shaped case.888Some readers may be troubled by the presumption that there could be experiments at all in CFR. But here we are on well-trodden ground. A community familiar with Boltzmann Brains should not be troubled by a few simple pieces of quantum optics. Such things will arise by chance in a sufficiently expansive CFR, and we may discuss their characteristic statistics.